OpenStudy (marigirl):

4e^(0.13t)-0.2e^(0.26t)=7.2 Please help me solve this.

OpenStudy (michele_laino):

hint: if I make this variable change: \[\Large {e^{0.13t}} = z\] I can rewrite the equation like below: \[\Large 4z - \frac{1}{5}{z^2} = \frac{{36}}{5}\] please solve that equation for \(z\)

OpenStudy (marigirl):

Appreciate it. Thanks!